OpenStudy (anonymous):

Suppose Mary is in consumer equilibrium. The marginal utility of good A is 30, and the price of good A is $2. a. if the price of good B is $4, the price of good C is $3, the price of good D is $1, and the price of all other goods and services is $5, what is the marginal utility of each goods mary is purchasing? b. If mary has chosen to keep $10 in savings, what is the ratio of MU to P for savings?

OpenStudy (anonymous):

If a consumer maximizes, MRS = price ratio. Thus, MRS ( AB) = MU(A)/MU(B) = 2/4 = 1/2 Thus, MU (b) = 60 MU for saving = 150. MU/P = 15.